### v2

# Positional Encoding in Transformer-Based Time Series Models: A Survey

Abstract
Recent advancements in transformer-based models have greatly improved time series analysis, providing robust solutions for tasks such as forecasting, anomaly detection, and classification. A crucial element of these models is positional encoding, which allows transformers to capture the intrinsic sequential nature of time series data. This survey systematically examines existing techniques for positional encoding in transformer-based time series models. We investigate a variety of methods, including fixed, learnable, relative, and hybrid approaches, and evaluate their effectiveness in different time series classification tasks. Our findings indicate that data characteristics like sequence length, signal complexity, and dimensionality significantly influence method effectiveness. Advanced positional encoding methods exhibit performance gains in terms of prediction accuracy, however, they come at the cost of increased computational complexity. Furthermore, we outline key challenges and suggest potential research directions to enhance positional encoding strategies. By delivering a comprehensive overview and quantitative benchmarking, this survey intends to assist researchers and practitioners in selecting and designing effective positional encoding methods for transformer-based time series models. The source code for the methods and experiments discussed in this survey is available on GitHub.
keywords:
Positional Encoding, Self-Attention, Transformer, Time Series.
1 Introduction
Time series analysis, which involves studying data collected over sequential time intervals, is fundamental across diverse domains including finance, healthcare, environmental monitoring, and industrial systems [1, 2]. The sequential nature of temporal data presents unique modeling challenges that distinguish it from other machine learning tasks, requiring specialized approaches to capture temporal dependencies, seasonal patterns, and long-range relationships effectively.
Traditional statistical methods such as Autoregressive Integrated Moving Average (ARIMA) models have long served as the foundation for time series analysis [1]. While these approaches provide interpretable results and work well for stationary data with clear patterns, they face significant limitations when dealing with non-linear relationships, high-dimensional multivariate data, and complex temporal dynamics that characterize modern time series applications [2].
The emergence of deep learning has revolutionized time series modeling by enabling the capture of complex non-linear patterns and long-range dependencies. Recurrent Neural Networks (RNNs) and their advanced variants, including Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU), became the dominant paradigm for sequential modeling [3, 4]. These architectures excel at processing sequential data through their inherent temporal structure, naturally handling variable-length sequences and maintaining memory of previous time steps. However, RNNs suffer from fundamental limitations including vanishing gradients, sequential processing bottlenecks, and difficulty capturing very long-range dependencies [5].
Convolutional Neural Networks (CNNs) offer an alternative approach to time series modeling by applying convolutional filters across the temporal dimension [6, 7]. While CNNs excel at capturing local temporal patterns and benefit from parallel processing, their fixed receptive fields constrain their ability to model long-term dependencies, and they require deep architectures or dilated convolutions to capture global patterns [8, 9].
The transformer architecture, introduced by Vaswani et al. [10], has fundamentally transformed sequence modeling by replacing recurrent connections with self-attention mechanisms. This paradigm shift enables parallel processing of entire sequences while maintaining the ability to capture long-range dependencies through direct attention computation between all sequence positions [11]. The unprecedented success of transformers in natural language processing, demonstrated by models like BERT [12] and GPT [13], has naturally motivated their application to time series analysis [14, 15, 16] with specialized architectures demonstrating superior performance across forecasting [17, 18, 19], classification [20, 21], and anomaly detection tasks [22, 23].
The adaptation of transformers to time series has revealed both remarkable opportunities and unique challenges. Unlike discrete text tokens with semantic relationships, time series involve continuous values with complex temporal dependencies, irregular sampling patterns, and domain-specific characteristics such as seasonality, trends, and cyclical behavior [24, 25]. The inherent permutation invariance of transformer architectures, while advantageous for certain tasks, becomes problematic for time series where temporal ordering is fundamental to the data’s meaning and predictive power [26].
This challenge has elevated positional encoding from a technical implementation detail to a critical design choice that fundamentally determines model effectiveness. Recent comprehensive analyzes have demonstrated that positional encoding methods significantly impact a model’s ability to handle sequence length extrapolation, temporal pattern recognition, and cross-domain generalization [27, 28]. The unique characteristics of temporal data—including non-stationarity, multi-scale patterns, and irregular sampling—require specialized positional encoding approaches that differ substantially from those developed for natural language processing [29].
Despite extensive research in transformer architectures, the time series community lacks a comprehensive evaluation of positional encoding methods specifically designed for temporal data characteristics. Although there are general transformer surveys [30] and NLP-focused positional encoding reviews are available [27], no systematic study has examined how these methods perform in various time series tasks and architectures. This gap is particularly critical given the unique properties of temporal data, including nonstationarity, seasonality, and varying sampling frequencies [24].
This survey addresses this critical gap by providing the first comprehensive empirical evaluation of positional encoding methods specifically for time series applications. We systematically analyze eight distinct positional encoding approaches across eleven diverse datasets representing different domains, sequence characteristics, and temporal patterns. Our evaluation spans two major transformer architectures to ensure the generalizability of findings and provides practical insights for method selection based on specific application requirements.
Our key contributions include:
  • A systematic comparison of eight positional encoding methods, along with their complexity analysis
  • An evaluation of these methods across diverse time series domains and different transformer architectures
  • The identification of performance patterns related to sequence characteristics such as length and dimensionality
  • practical guidelines for method selection
Results demonstrate that advanced methods like Stochastic Positional Encoding (SPE) and Transformer with Untied Positional Encoding (TUPE) [31] consistently outperform traditional approaches, with effectiveness varying significantly by sequence length, dimensionality, and application domain.
The remainder of this survey is organized as follows: Section 2 provides a comprehensive literature review of transformer-based time series models and positional encoding methods. Section 3 presents background on self attention mechanism positional encoding fundamentals. Section 4 details our systematic analysis of different positional encoding categories, and Sections 5-6 discuss evaluation methodologies and empirical results.
2 Literature Review
2.1 Attention-based models
The evolution of attention mechanisms in time series analysis has progressed through several distinct phases, beginning with early attention-augmented architectures and culminating in sophisticated transformer-based models specifically designed for temporal data. Before full transformer architectures gained prominence, researchers explored attention mechanisms as enhancements to existing sequential models, building upon foundational attention work in neural machine translation [32, 33]. Early work focused on augmenting traditional neural networks with attention capabilities to improve their ability to model temporal dependencies and handle variable-length sequences.
The Dual-Stage Attention-Based RNN (DA-RNN) [34] pioneered the use of attention at both input and temporal dimensions, allowing models to selectively focus on relevant features and time steps. This approach demonstrated significant improvements in financial time series prediction by adaptively extracting relevant dependencies, establishing the foundation for attention-based temporal modeling. Similar dual attention strategies were subsequently developed for multivariate forecasting [35] and clinical time series analysis [36].
Attention-based encoder-decoder architectures adapted sequence-to-sequence learning paradigms for time series forecasting [37]. These models showed that attention mechanisms could effectively handle variable-length sequences and improve long-term prediction accuracy by focusing on relevant historical information. This work demonstrated that attention could address some of the fundamental limitations of purely sequential processing approaches. The success of self-attention motivated the development of specialized variants designed for temporal data characteristics. Cross Attention Stabilized Fully Convolutional Neural Network (CA-SFCN) [38] combined convolutional layers with two types of self-attention: temporal attention (TA) for identifying important time steps and variable attention (VA) for capturing relationships between different variables in multivariate time series. This dual attention approach showed that different attention mechanisms could be specialized for different aspects of temporal modeling. Gated Transformer Networks (GTN) [39] introduced two-tower multi-headed attention architectures that capture discriminative information from input series, using learnable gating mechanisms to balance between different attention patterns. This approach demonstrated that attention mechanisms could be made more adaptive to specific temporal patterns, paving the way for more sophisticated attention-based architectures. The computational challenges of applying attention to long time series motivated the development of efficient attention mechanisms. Sparse attention patterns [40] reduce computational complexity by focusing on subsets of time steps, while maintaining the ability to capture relevant dependencies. These efficiency improvements were crucial for making attention-based approaches practical for real-world time series applications with long sequences. Complementary efficiency strategies include locality-enhanced transformers [11] and reformer architectures [41] that reduce memory bottlenecks in long sequence processing.
Recent developments have focused on adaptive attention mechanisms that adjust based on temporal context [42, 43, 44].
2.2 Transformer-Based Models
Building upon the foundation of attention mechanisms, transformer architectures represent a paradigm shift towards pure attention-based sequence modeling. In time series analysis, adaptations of Transformers have led to specialized architectures tailored to exploit their strengths while addressing temporal-specific challenges. The Transformer-based framework for multivariate time series representation learning [20] integrates batch normalization to stabilize training, enhancing model performance on complex multivariate datasets. This framework underscores the significance of normalization techniques in fortifying Transformer models for diverse time series tasks. The multi-head attention mechanism has been extensively adapted for time series applications, with different heads specializing in various temporal patterns. The success of multi-head attention in transformers inspired specialized variants for temporal data. The Temporal Fusion Transformer (TFT) [45] introduced specialized attention heads for different temporal patterns: some heads focus on short-term dependencies, while others capture long-term trends. The model also incorporates variable selection networks to handle high-dimensional input features common in time series applications. Contemporary approaches have explored hierarchical and multi-scale positional encodings [46, 47, 48].
Moreover, inspired by vision Transformers, patch embeddings have been adapted for time series data. This approach, similar to the Vision Transformer (ViT) paradigm [49], segments time series into patches for localized processing. By encapsulating both local and global patterns, patch embeddings facilitate hierarchical pattern recognition, thereby improving model interpretability and performance in time series analysis.
However, transformers face unique challenges in time series applications. Their inherent permutation invariance necessitates explicit positional information to distinguish temporal ordering [15, 50, 20]. Additionally, the global attention mechanism may dilute focus on local patterns, and the absence of natural recency bias requires architectural adaptations for optimal time series performance.
2.3 Positional Encoding in Time Series
The adaptation of positional encoding methods to time series has revealed fundamental differences between temporal and linguistic data that require specialized approaches.
Positional encoding (PE) mechanisms are therefore crucial for enabling transformers to learn temporal relationships effectively, as self-attention mechanisms are inherently permutation-invariant. [51]. Vaswani et al. [10] initially introduced fixed sinusoidal encodings to embed positional information for temporal differentiation. While these provided a foundational approach, they can struggle with varying sequence lengths and dynamic temporal relationships common in time series data [52, 45]. Subsequently, Shaw et al. [52] advanced the field with learnable and relative positional encodings, enhancing temporal relationship modeling capabilities.
Recent innovations have specifically targeted time series challenges. Foumani et al. [21] developed Time Absolute Position Encoding (tAPE) and Efficient Relative Position Encoding (eRPE) for enhanced temporal dependency modeling, alongside the ConvTran model—a hybrid architecture fusing convolutional layers with transformer mechanisms. Other advanced methods include Transformer with Untied Positional Encoding (TUPE) [31] and Convolutional Stochastic Position Encoding (ConvSPE) [53], which address content-position decoupling and linear complexity maintenance respectively. Hybrid approaches, combining absolute and relative strategies, have also been proposed to leverage complementary strengths for comprehensive temporal information representation, aiming to address some of the inherent limitations of transformer architectures in time series analysis.
Recent work has systematically evaluated PE methods for time series applications [54, 55]. Advanced techniques include frequency-based encodings [56, 57], learnable approaches [58, 59], and hybrid strategies [60, 61]. Specialized methods for irregular time series [62, 63] and long sequences [64, 65] have also been developed.
Table 1: Summary of Positional Encoding Techniques. Abs=Absolute, Rel=Relative, Hyb=Hybrid, Add=Additive, Att=Attention, Comb=Combined, F=Fixed, L=Learnable, M=Mixed.
Method Type Inject. Learn. Params Memory Complex.
Sin. PE [10] Abs Add F 0 $\mathcal{O}(Ld)$ $\mathcal{O}(Ld)$
Learn. PE [10] Abs Add L LdLd $\mathcal{O}(Ld)$ $\mathcal{O}(Ld)$
RPE [52] Rel Att F (2L1)dl(2L{-}1)dl $\mathcal{O}(L^{2}d)$ $\mathcal{O}(L^{2}d)$
tAPE [21] Abs Add F LdLd $\mathcal{O}(Ld)$ $\mathcal{O}(Ld)$
RoPE [66] Hyb Att F 0 $\mathcal{O}(Ld)$ $\mathcal{O}(L^{2}d)$
eRPE [21] Rel Att L 2L12L-1 $\mathcal{O}(L^{2}+L)$ $\mathcal{O}(L^{2})$
TUPE [31] Hyb Att L $2dl$ $\mathcal{O}(Ld{+}d^{2})$ $\mathcal{O}(Ld{+}d^{2})$
ConvSPE [53] Rel Att L 3Kdh+dl3Kdh{+}dl $\mathcal{O}(LKR)$ $\mathcal{O}(LKR)$
T-PE [29] Hyb Comb M $2d^{2}l/h{+}(2L{+}2l)d$ $\mathcal{O}(L^{2}d)$ $\mathcal{O}(L^{2}d)$
ALiBi [67] Rel Att F 0 $\mathcal{O}(L^{2}h)$ $\mathcal{O}(L^{2}h)$
Table 1 provides a comprehensive taxonomy of the positional encoding techniques discussed, based on criteria such as technique type, injection method, learning strategy, parameters, and time complexity. The parameter counts include all learnable parameters across layers, where $l$ represents the number of transformer layers, $h$ is the number of attention heads, and $d$ is the hidden dimension. For methods using relative positioning (RPE, eRPE), we include the full attention matrix parameters. For ConvSPE, K represents the convolutional kernel size that controls the local neighborhood for convolution operations.
3 Positional Encoding Analysis
This section offers a fundamental explanation of self-attention and a summary of existing positional encoding models. It’s important to distinguish between positional encoding, which refers to the approach used to incorporate positional information (e.g., absolute or relative), and positional embedding, which denotes the numerical vector associated with that encoding.
3.1 Problem definition and notation
Given a time series dataset $X$ consisting of $n$ samples, $X=\{x_{1},x_{2},\dots,x_{n}\}$, where each sample $x_{i}=\{x_{i1},x_{i2},\dots,x_{iL}\}$ represents a $d_{x}$-dimensional time series of length $L$, such that $x_{i}\in\mathbb{R}^{L\times d_{x}}$. Additionally, let $Y=\{y_{1},y_{2},\dots,y_{n}\}$ denote the corresponding set of labels, where $y_{t}\in\{1,2,\dots,c\}$ and $c$ is the number of classes. The objective is to train a neural network classifier to map the input dataset $X$ to its corresponding label set $Y$.
3.2 Self-Attention Mechanism
Figure 1: Scaled dot-product attention (left) and multi-head attention (right) mechanisms in transformer architecture.
The self-attention mechanism is a key component of Transformer models, enabling them to capture long-range dependencies in sequential data. Given an input sequence $X\in\mathbb{R}^{L\times d_{x}}$, the self-attention mechanism computes a weighted representation of each element in the sequence by attending to all other elements, as illustrated in Fig. 1.
First, three linear projections are applied to the input sequence to obtain the query ($Q$), key ($K$), and value ($V$) matrices:
$Q=XW_{Q},\quad K=XW_{K},\quad V=XW_{V},$
followed by scaled dot-product attention:
$Attention(Q,K,V)=softmax\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$
The output of the self-attention mechanism is a weighted sum of the value vectors, where the weights are determined by the attention scores. This process allows the model to focus on the most relevant parts of the sequence when generating each element of the output.
The theoretical foundations of attention mechanisms have been extensively analyzed [68, 69], revealing their universal approximation capabilities for sequence-to-sequence functions. Recent work has provided deeper insights into attention patterns and their interpretability [70], which is particularly relevant for understanding how different positional encodings influence attention weight distributions in time series contexts.
Multi-head attention concatenates parallel attention heads to learn diverse representations:
$\text{MultiHead}(Q,K,V)$ $=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h})W_{O}$
Transformers lack an inherent notion of sequence order, as self-attention operates on sets of inputs without considering their positions. To address this, positional encoding is introduced to incorporate positional information into the model, allowing it to understand the order of elements in a sequence.
Let the input sequence be $X\in\mathbb{R}^{L\times d_{x}}$, where $L$ is the sequence length and $d_{x}$ is the feature dimension. Positional encoding adds a unique vector to each position in the sequence, resulting in a new representation $X^{\prime}$:
$X^{\prime}=X+P,$
where $P\in\mathbb{R}^{L\times d_{x}}$ is the positional encoding matrix.
The theoretical foundations of positional encoding in sequential models have been extensively analyzed from both signal processing [71, 72] and cybernetic control perspectives [73, 74].
3.3 Absolute Positional Encoding
Absolute positional encoding assigns unique vectors to each temporal position. The sinusoidal approach [10] provides computational efficiency by using predetermined trigonometric functions, eliminating the need for learned positional representations.
3.3.1 Fixed Absolute Positional Encoding
The fixed absolute positional encoding is defined as follows:
$PE_{(pos,2i)}$ $=\sin\left(\frac{pos}{10000^{\frac{2i}{d}}}\right)$
$PE_{(pos,2i+1)}$ $=\cos\left(\frac{pos}{10000^{\frac{2i}{d}}}\right)$
where pospos is the position (time step) in the sequence, $i$ is the dimension, and $d$ is the dimensionality of the positional encoding.
The use of sinusoidal functions ensures that the encodings for nearby positions are similar, which can help the model learn relationships between closely spaced time steps. However, this approach has limitations when dealing with sequences of varying lengths or non-uniform time intervals, which are common in time series data.
3.3.2 Learnable Positional Encoding
Learnable positional encoding provides more flexibility by allowing the model to learn the positional representations during training. This can be advantageous for time series tasks where the relationships between time steps may be complex or nonlinear. The learnable positional encoding is typically defined as:
$PE_{learnable}({pos})=W_{{pos}}$
where $W_{pos}$ is an $d_{\text{model}}$-dimensional vector representing the learnable positional encoding parameters for position pospos.
During the forward pass of the neural network, the learnable positional encoding is added to the input embeddings $x$:
$x=x+PE_{learnable}[:x.size(1),:]$
where $x$ represents the input tensor at a given position and $x.\text{size}(1)$ denotes the length of the sequence up to the current position. This addition enables the model to learn position-specific transformations that can capture dependencies and patterns within sequences.
The flexibility and adaptability of learnable positional encoding make it a powerful tool for enhancing the modeling of temporal dependencies and patterns in time series data within transformer architectures.
Then, the self-attention mechanism is modified as follows:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{(Q+PE)\cdot(K+PE)^{T}}{\sqrt{d_{k}}}\right)V$
where $Q$, $K$, and $V$ are the query, key, and value matrices, respectively, $PE$ represents the positional encoding matrix, and $d_{k}$ is the dimensionality of the key vectors [10].
3.4 Relative Positional Encoding
Shaw et al. [52] proposed modifying the attention computation to include relative position information:
$e_{ij}=\frac{(x_{i}W_{Q})(x_{j}W_{K}+a^{K}_{ij})^{T}}{\sqrt{d_{z}}}$
$e_{ij}=\frac{(x_{i}W_{Q})(x_{j}W_{K})^{T}+(x_{i}W_{Q})(a^{K}_{ij})^{T}}{\sqrt{d_{z}}}$
$a^{K}_{ij}=W_{K}^{r}r_{i-j}$
where $x_{i}$ and $x_{j}$ are the input representations at positions $i$ and $j$, $W_{Q}$ and $W_{K}$ are the learnable projection matrices for queries and keys, and $a^{K}_{ij}$ is the relative positional encoding defined as $a^{K}_{ij}=W_{K}^{r}r_{i-j}$, where $r_{i-j}$ is the learnable relative positional embedding for the distance iji-j.
This formulation allows the model to better capture temporal dependencies that depend on the relative position between time steps, which is particularly useful for tasks such as time series forecasting or classification when dealing with varying sequence lengths [52, 75].
Figure 2: Absolute vs. relative positional encoding: absolute methods add position vectors to input embeddings, while relative methods directly incorporate positional relationships in attention computation.
As illustrated in Fig. 3, positional encoding methods can be combined to leverage the advantages of both absolute and relative approaches.
3.5 Time Absolute Position Encoding (tAPE)
Foumani et al. [21] identified that traditional sinusoidal encoding loses distance awareness in lower-dimensional embeddings. tAPE addresses this by incorporating sequence length into frequency computation:
$\omega_{k}^{new}=k\times\frac{d_{model}}{L}$
$PE_{pos,2i}=\sin(pos\times\omega_{k}{new})$
$PE_{pos,2i+1}=\cos(pos\times\omega_{k}{new})$
This modification ensures consistent distance awareness across different embedding dimensions and sequence lengths, making it particularly suitable for diverse time series applications.
3.6 Rotary Positional Encoding (RoPE)
Rotary Positional Encoding (RoPE) [66] uses rotation matrices to encode absolute positions while incorporating explicit relative position dependencies. This method applies rotations to query and key representations based on their absolute positions before computing dot-product attention:
$e_{ij}=\frac{(x_{i}W^{Q}+R_{n\theta})(x_{j}W^{K}+R_{n\theta})^{T}}{\sqrt{d_{z}}}$ (1)
where $R_{n\theta}$ is a rotation matrix that rotates by $n\theta$ radians:
$R_{n\theta}=\begin{bmatrix}\cos(n\theta)&-\sin(n\theta)\\ \sin(n\theta)&\cos(n\theta)\end{bmatrix}$ (2)
where $n$ is the position of the data, and $\theta$ is a scalar value (angle) chosen to rotate the value. RoPE combines the benefits of both absolute and relative PE, ensuring that attention calculations depend on relative distances while maintaining absolute position information. This hybrid approach has shown particular effectiveness in capturing long-range dependencies in sequential data.
3.7 Efficient Relative Position Encoding (eRPE)
This method enhances traditional relative positioning by applying bias terms after the softmax operation:
$\alpha_{i}=\sum_{j\in L}\left(\underbrace{\frac{\exp(e_{i,j})}{\sum_{k\in L}\exp(e_{i,k})}}{A{i,j}}+w_{i-j}\right)x_{j}$
where:
where $w\in\mathbb{R}^{2L-1}$ represents learnable relative position biases. The post-softmax application sharpens positional distinctions, making the method particularly effective for tasks requiring precise temporal localization.
3.8 Transformer with Untied Positional Encoding (TUPE)
The Transformer with Untied Positional Encoding (TUPE) [31] introduces a novel approach by explicitly separating content and positional information in the attention computation, unlike traditional methods that simply add positional embeddings to input embeddings. The attention score in TUPE is computed as:
$\alpha_{ij}=\frac{1}{\sqrt{2d}}(x_{i}^{l}W^{Q,l})\cdot(x_{j}^{l}W^{K,l})^{T}+\text{reset}_{\theta}(\beta_{ij},i,j)$
where $\beta_{ij}=\frac{1}{\sqrt{2d}}(p_{i}U^{Q})(p_{j}U^{K})^{T}$ represents position-to-position interactions. This separation prevents interference between content and positional patterns, allowing each component to be learned independently. For time series data, this is particularly advantageous as it enables the model to capture both temporal patterns (through positional correlations) and feature relationships (through content correlations) without mutual interference.
3.9 Convolutional Sinusoidal Positional Encoding (ConvSPE)
Convolutional Stochastic Positional Encoding (convSPE) [53] enables relative positional encoding through a convolution-based approach that maintains linear complexity. The method begins by sampling a random matrix $Z_{d}\in\mathbb{R}^{M\times R}$ with i.i.d. standard Gaussian entries, then applies learnable convolutional filters $\Phi^{Q}$ and $\Phi^{K}$ to obtain query and key encodings:
$Q_{d}=Z_{d}*\Phi^{Q},\quad K_{d}=Z_{d}*\Phi^{K}$
where * denotes convolution. The convolution operation naturally induces relative position awareness through the sliding window mechanism. The theoretical foundation ensures that for large $R$, the cross-covariance structure approximates the desired positional attention kernel:
$P_{d}\approx\frac{\bar{Q}_{d}\bar{K}_{d}^{T}}{R}$ (3)
This formulation maintains linear complexity while effectively capturing relative positions through the convolutional structure, making it particularly suitable for long time series sequences.
3.10 Temporal Positional Encoding (T-PE)
This method combines geometric and semantic components to capture both positional and content-based temporal relationships: The semantic component adjusts attention scores by incorporating token similarity between positions $i$ and $j$:
$T\text{-}PE(i)=PE(i)+S(i,j)$
$S(i,j)=\exp\left(-\frac{\|x_{i}-x_{j}\|^{2}}{2\sigma^{2}}\right)$
The geometric component maintains traditional positional information, while the semantic component $S(i,j)$ measures similarity between tokens at different positions, enabling the model to identify recurring patterns regardless of their absolute positions.
3.11 Attention with Linear Biases (ALiBi)
This method  [67] introduces a fundamentally different approach by eliminating traditional positional embeddings. Instead, ALiBi directly modifies attention scores by adding linear biases proportional to the distance between query and key positions:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+\mathbf{m}\cdot\mathbf{B}\right)V$
where $\mathbf{B}$ is the bias matrix with elements $B_{ij}=-|i-j|$ representing the absolute distance between positions $i$ and $j$, and $\mathbf{m}$ is a head-specific slope parameter controlling bias strength. The slope values are computed using a geometric sequence:
$m_{h}=2^{-\frac{8h}{H}}\text{ for }h=1,2,\ldots,H$
For time series applications, ALiBi uses symmetric distance calculations $B_{ij}=-|i-j|$ to enable bidirectional attention across temporal sequences. This parameter-free approach naturally encodes local temporal pattern preference while permitting long-range dependencies and enables extrapolation to longer sequences than seen during training.
Figure 3: The left matrix represents standard attention score computation $q_{i}\cdot k_{j}$ between queries and keys. The right matrix shows ALiBi’s linear bias values that are added to attention scores before softmax normalization.
4 Experimental Setup
The experimental setup involves comparing different positional encoding methods in two distinct Transformer-based architectures for time series data. The Time Series Transformer [20], shown in Figure 4(a), processes the raw multivariate time series data directly. For a time series with $L$ timesteps and $d$ channels, it treats each timestep as a token, resulting in $L$ tokens of dimension $d$ that are fed into the transformer layers. This architecture incorporates normalization layers after both the attention and feed-forward components to stabilize training and address internal covariate shifts. In contrast, the patch embedding transformer [49], illustrated in Figure 4(b), first transforms the input data using convolutional operations to create overlapping patches. Each patch spans multiple timesteps but maintains the full channel dimension, effectively reducing the sequence length while enriching the feature representation of each token. This results in $L/p$ tokens (where $p$ is the patch size) with increased feature dimensionality before being processed by the transformer layers. These architectures were chosen to evaluate positional encoding methods under different sequence processing paradigms: direct temporal processing in the Time Series Transformer versus hierarchical feature extraction through patch embedding.
(a) Time Series Transformer
(b) Patch Embedding Transformer
Figure 4: Transformer architectures for time series classification.
4.1 System Configuration
We conducted all experiments on a high-performance Linux server with the following specifications: an AMD EPYC 7513 32-Core Processor (2 sockets, 128 threads), one NVIDIA RTX A5000 GPUs (24GB GDDR6 each, CUDA version 12.2, Driver version 535.113.01), 503GB of RAM, and a dual storage system consisting of an 892.7GB NVMe SSD and a 10.5TB HDD. The operating system used was Ubuntu 20.04.6 LTS (Kernel 5.4.0-200-generic). All implementations were carried out using Python 3.8.10 and PyTorch 2.4.1+cu121.
4.2 Datasets
A comprehensive set of time series datasets from various domains is employed to evaluate the positional encoding methods’ effectiveness across diverse applications. The selected datasets include:
  • Sleep: EEG (Electroencephalogram) signals with 178 time steps per instance. The dataset contains 478,785 training instances and 90,315 test instances, classified into five different stages of sleep.
  • ElectricDevices: Electrical consumption data of household devices with 96 time steps. It contains 8,926 training instances and 7,711 test instances across seven device classes [76].
  • FaceDetection: EEG signals collected during face detection tasks, with 62 time steps per instance. The dataset has 5,890 training instances and 3,524 test instances for a binary classification task.
  • MelbournePedestrian: Traffic data capturing pedestrian flow in Melbourne. Each instance contains 24 time steps, with a 10-class classification task, comprising 1,194 training instances and 2,439 test instances.
  • SharePriceIncrease: Financial time-series data with 60 time steps per instance, used to predict whether a company’s share price will increase. The dataset has 965 training instances and 965 test instances, with two classes.
  • LSST: Astronomical time-series data from the Large Synoptic Survey Telescope, containing 36 time steps per instance. The dataset includes 2,459 training instances and 2,466 test instances for a 14-class classification task of celestial objects.
  • RacketSports: Human activity recognition data from racket sports like tennis and badminton. Each instance contains 30 time steps, with 151 training instances and 152 test instances across four activity classes.
  • SelfRegulationSCP1: EEG data related to self-regulation through slow cortical potentials (SCPs), consisting of 896 time steps per instance. The dataset contains 268 training instances and 293 test instances for binary classification.
  • UniMiB-SHAR: Sensor data for human activity recognition, with 151 time steps per instance. The dataset comprises 4,601 training instances, 1,454 validation instances, and 1,524 test instances across nine activity classes [77].
  • RoomOccupancy (Sensors): Time-series data from non-intrusive environmental sensors, such as temperature, light, sound, and CO2, used to estimate the precise number of occupants in a room. The dataset has 10,129 instances with 18 features and is classified into four classes [78].
  • EMGGestures (EMG): Electromyographic (EMG) signals recorded using the MYO Thalmic bracelet during various gesture executions. Each instance consists of 30 time steps with 9 features per time step. The dataset is classified into eight distinct gesture classes. The dataset is primarily used for human activity recognition in health and medicine-related applications [79].
    The diversity of our dataset collection reflects the heterogeneous nature of time series applications across domains. Our selection criteria prioritize datasets that have been standardized in the time series classification literature [80, 81], ensuring reproducibility and enabling comparison with established baselines. Additionally, we include datasets from the Monash Time Series Forecasting Archive [82] to represent real-world forecasting scenarios, though our primary evaluation focuses on classification tasks for the reasons outlined in Section 1.
    The characteristics of these datasets are summarized in Table 2, which provides a comprehensive overview of their key properties including training and test set sizes, sequence lengths, number of classes, and channels.
Table 2: Characteristics of time series datasets used in the experiments.
Dataset Train Size Test Size Length Classes Channel Type
Sleep 478,785 90,315 178 5 1 EEG
ElectricDevices 8,926 7,711 96 7 1 Device
FaceDetection 5,890 3,524 62 2 144 EEG
MelbournePedestrian 1,194 2,439 24 10 1 Traffic
SharePriceIncrease 965 965 60 2 1 Financial
LSST 2,459 2,466 36 14 6 Other
RacketSports 151 152 30 4 6 HAR
SelfRegulationSCP1 268 293 896 2 6 EEG
UniMiB-SHAR 4,601 1,524 151 9 3 HAR
RoomOccupancy 8,103 2,026 30 4 18 Sensor
EMGGestures 1,800 450 30 8 9 EMG
4.3 Evaluation Procedure
Our evaluation methodology draws inspiration from comprehensive benchmarking practices established in both the NLP transformer literature [13] and time series analysis communities [16]. We employ stratified sampling and cross-validation techniques to ensure robust statistical evaluation, following guidelines established for deep learning evaluation in temporal domains [83]. The experiments are designed to assess how well each positional encoding technique captures temporal dependencies and patterns in the data and how it affects the overall model performance on each architecture. The evaluation procedure includes:
  1. 1.
    Model Training: Each model is trained on the training portion of the datasets using standard optimization techniques and hyperparameters.
  2. 2.
    Model Evaluation: The trained models are evaluated on the validation and test sets to measure their performance according to the specified metrics.
  3. 3.
    Comparative Analysis: The results are compared across different models to determine the effectiveness of each positional encoding method.
The experiments aim to provide insights into the strengths and weaknesses of each positional encoding technique and identify the scenarios where each method performs best.
5 Results
Our experimental evaluation encompasses eight distinct positional encoding methods tested across eleven diverse time series datasets using two transformer architectures. The methods range from traditional approaches (PE, LPE) to relative encodings (RPE, eRPE) and advanced techniques (T-PE, SPE, TUPE), evaluated on both time series transformer and transformer with patch embedding architectures. The classification accuracies are presented in Tables 3 and 4, with performance visualizations provided through heatmaps in Figures 6(b) and 6(a).
5.1 Overall Performance
The experimental results demonstrate a consistent superiority of advanced positional encoding methods across both architectural configurations. SPE achieves the highest average ranking (1.727 for time series transformer, 2.090 for patch embedding), closely followed by TUPE (1.909 and 2.272) and T-PE (2.636 and 2.363). This performance hierarchy remains remarkably stable across both architectures, though the patch embedding architecture exhibits more balanced performance among the top-performing methods, as shown in Figures 5(a) and 5(b). These findings align with recent studies showing that advanced positional encodings provide significant benefits in complex temporal modeling tasks [84, 85].
5.2 Sequence Length Analysis
The effectiveness of positional encoding methods shows a strong correlation with sequence length. For sequences exceeding 100 time steps, such as SelfRegulationSCP1 (896 steps) and Sleep (178 steps), advanced methods demonstrate substantial improvements, achieving 5-6% higher accuracy compared to fixed encoding approaches. Medium-length sequences (50-100 steps) show moderate improvements of 3-4%, while shorter sequences (less than 50 steps) exhibit smaller but consistent gains of 2-3%. The correlation between sequence length and encoding effectiveness has been observed in other domains [86, 87], confirming the importance of sophisticated positioning strategies for extended sequences.
5.3 Dataset-Specific Performance
5.3.1 Biomedical Signals
In EEG datasets (Sleep, FaceDetection, SelfRegulationSCP1), TUPE demonstrates optimal performance, achieving accuracies of 88.8%, 70.5%, and 90.7% respectively under the time series transformer architecture. SPE shows comparable performance levels, particularly excelling in EMG data with 77.1% accuracy. Both methods demonstrate superior capability in capturing the long-range dependencies characteristic of biomedical signals.
5.3.2 Environmental and Sensor Data
For sensor-based datasets (ElectricDevices, RoomOccupancy), SPE exhibits superior performance (75.5% and 95.1% respectively). While TUPE maintains competitive accuracy, relative encoding methods show improved performance in local pattern recognition. The performance advantage of advanced methods is particularly pronounced in datasets characterized by complex temporal dependencies.
5.4 Architectural Impact
While both architectures maintain similar method rankings, their performance characteristics exhibit notable differences. The time series transformer architecture shows more distinct performance gaps between method tiers, while the patch embedding architecture demonstrates closer competition among advanced methods. This suggests that architectural choice significantly influences the relative advantages of different encoding strategies, though the superiority of advanced methods remains consistent.
5.5 Dimensionality and Class Complexity
Performance patterns vary significantly with input dimensionality and class complexity. Higher-dimensional datasets (more than 5 channels) show more consistent improvements across methods, while lower-dimensional data exhibits more variable benefits. Classification tasks with more classes (exceeding 7) demonstrate smaller performance gaps between methods, though the ranking hierarchy remains stable.
The comprehensive results establish the robust superiority of advanced positional encoding methods, particularly in complex time series tasks with longer sequences or higher dimensionality. The consistent performance patterns across different architectures and dataset characteristics suggest the broader applicability of these advanced encoding strategies in time series analysis.
Table 3: Accuracy of Positional Encoding Methods in Time Series Transformer
Dataset Absolute PE Relative PE Hybrid PE
Sin. PE Learn. PE tAPE RPE eRPE SPE ALiBi RoPE T-PE TUPE
Sleep (EEG) 0.854 0.862 0.871 0.867 0.875 0.883 0.861 0.864 0.881 0.888
ElectricDevices 0.682 0.697 0.718 0.723 0.731 0.755 0.702 0.693 0.741 0.743
FaceDetection 0.627 0.654 0.633 0.672 0.686 0.698 0.638 0.647 0.694 0.705
MelbournePedestrian 0.675 0.704 0.701 0.729 0.742 0.754 0.688 0.695 0.748 0.762
SharePriceIncrease 0.725 0.748 0.744 0.739 0.772 0.785 0.733 0.732 0.779 0.778
LSST 0.581 0.596 0.609 0.613 0.632 0.628 0.601 0.605 0.629 0.623
RacketSports 0.742 0.762 0.761 0.775 0.804 0.813 0.737 0.764 0.811 0.819
SelfRegulationSCP1 0.849 0.857 0.853 0.871 0.891 0.901 0.864 0.842 0.898 0.907
UniMib 0.841 0.855 0.852 0.865 0.876 0.879 0.853 0.859 0.882 0.878
RoomOccupancy 0.912 0.922 0.928 0.931 0.942 0.951 0.925 0.922 0.947 0.948
EMGGestures 0.713 0.709 0.721 0.735 0.751 0.771 0.705 0.715 0.763 0.766
Avg Rank 9.636 6.454 6.363 5.272 3.727 1.727 8.181 7.27 2.636 1.909
Table 4: Accuracy with Positional Encoding Methods in Transformer + Patch Embedding
Dataset Absolute PE Relative PE Hybrid PE
Sin. PE Learn. PE tAPE RPE eRPE SPE ALiBi RoPE T-PE TUPE
Sleep (EEG) 0.841 0.852 0.851 0.858 0.864 0.876 0.853 0.844 0.872 0.879
ElectricDevices 0.673 0.694 0.693 0.761 0.762 0.811 0.684 0.713 0.763 0.779
FaceDetection 0.618 0.642 0.653 0.671 0.678 0.674 0.626 0.635 0.686 0.695
MelbournePedestrian 0.668 0.702 0.682 0.724 0.733 0.753 0.672 0.690 0.742 0.745
SharePriceIncrease 0.715 0.742 0.735 0.754 0.774 0.785 0.721 0.731 0.781 0.774
LSST 0.573 0.582 0.584 0.596 0.611 0.601 0.592 0.583 0.602 0.595
RacketSports 0.733 0.756 0.761 0.767 0.801 0.775 0.722 0.749 0.792 0.805
SelfRegulationSCP1 0.836 0.844 0.842 0.829 0.856 0.883 0.849 0.831 0.871 0.875
UniMib 0.832 0.844 0.833 0.854 0.864 0.867 0.841 0.836 0.871 0.865
RoomOccupancy 0.905 0.911 0.922 0.918 0.929 0.934 0.915 0.917 0.931 0.937
EMGGestures 0.697 0.692 0.713 0.722 0.741 0.746 0.702 0.707 0.752 0.742
Avg Rank 9.636 6.454 6.363 5.272 3.363 2.090 7.818 7.363 2.363 2.272
(a) Patch Embedding Transformer
(b) Time Series Transformer
Figure 5: Average rank of various combinations of absolute and relative position encodings across different transformer architectures.
(a) Patch Embedding Transformer
(b) Time Series Transformer
Figure 6: Heatmaps illustrating the accuracy (%) of various positional encoding methods across different datasets. Each cell represents the accuracy for a specific dataset and positional encoding method, with darker colors indicating higher accuracy.
6 Discussion
6.1 Impact of Sequence Characteristics
Our comprehensive analysis reveals that the effectiveness of positional encoding methods is fundamentally influenced by sequence characteristics, particularly sequence length and temporal complexity. The superior performance of advanced methods (SPE, TUPE) becomes increasingly pronounced with sequence length, demonstrating up to 5.8% improvement in sequences exceeding 100 time steps, compared to modest gains of 2-3% in shorter sequences. This pattern strongly suggests that sophisticated encoding strategies are particularly crucial for applications involving extended temporal dependencies, such as EEG analysis and complex sensor data processing.
The relationship between sequence length and encoding method effectiveness can be attributed to two key factors. First, longer sequences present more opportunities for capturing complex temporal dependencies, allowing advanced methods to leverage their sophisticated encoding mechanisms more effectively. Second, the increased dimensionality of longer sequences provides a richer feature space for learning position-aware representations, enabling methods like TUPE and SPE to better differentiate between temporal patterns at various scales.
6.2 Architectural Considerations
The comparative analysis of transformer architectures reveals distinctive performance patterns that provide insights into the interaction between architectural choices and positional encoding strategies. While both architectures maintain similar method rankings, time series transformer architecture demonstrates more pronounced performance gaps between method tiers. This suggests that the normalization process amplifies the benefits of advanced encoding strategies, possibly by stabilizing the learning of complex positional patterns.
Conversely, the patch embedding architecture shows more balanced performance among top methods, particularly benefiting multivariate datasets where local feature interactions are crucial. This architectural characteristic suggests that patch embedding’s inherent ability to capture local patterns complements the global temporal dependencies encoded by advanced positional encoding methods, resulting in more uniform performance improvements across different encoding strategies.
6.3 Dataset-Specific Effectiveness
Dataset characteristics significantly influence encoding method performance, revealing important patterns for practical applications. In biomedical signals, particularly EEG datasets, TUPE and SPE consistently excel, suggesting these methods’ superior capability in capturing complex physiological patterns. Their effectiveness likely stems from their ability to model both local temporal dynamics and long-range dependencies characteristic of biological signals. The superior performance on biomedical signals reflects the methods’ ability to capture complex physiological feedback loops and temporal dependencies characteristic of biological systems [88, 89].
For environmental and sensor data, SPE’s superior performance (up to 95.1% accuracy) indicates its effectiveness in handling regular sampling patterns and clear temporal structures. This suggests that SPE’s encoding strategy is particularly well-suited for datasets with consistent temporal patterns and well-defined periodicities. The performance advantages in these domains demonstrate the importance of matching encoding strategies to specific data characteristics.
6.4 Computational Considerations
The practical deployment of positional encoding methods requires careful consideration of computational trade-offs beyond accuracy metrics [90]. We analyze the computational characteristics of different PE methods to provide guidance for resource-constrained deployment scenarios.
Table 5: Training Time Analysis and Efficiency Metrics for PE methods on Melbourne Pedestrian Dataset in Transformer + Patch Embedding (100 epochs)
Method Time (s) Ratio Accuracy
Sin. PE 48.2 1.00 0.668
Learn. PE 60.1 1.25 0.702
RPE 128.4 2.66 0.724
tAPE 54.0 1.12 0.682
RoPE 67.8 1.41 0.690
eRPE 142.8 2.96 0.733
TUPE 118.3 2.45 0.745
ConvSPE 101.6 2.11 0.753
T-PE 134.7 2.79 0.742
AliBi 93.8 1.94 0.672
Figure 7: Performance vs computational complexity trade-off on Melbourne dataset. ConvSPE achieves the optimal efficiency frontier, delivering highest accuracy with moderate computational overhead.
[…]
6.5 Practical Guidelines for Method Selection
Based on our empirical analysis, we provide a systematic decision framework for positional encoding selection in transformer-based time series models.
6.5.1 Sequence Length-Based Recommendations
Short sequences ($L\leq 50$):
Use Learnable PE or tAPE. Advanced methods provide minimal gains (2–3%) that rarely justify computational overhead.
Medium sequences ($50<L\leq 100$):
SPE or eRPE recommended for 3–4% accuracy improvements. Consider T-PE when interpretability is required.
Long sequences ($L>100$):
TUPE for complex patterns, SPE for regular data, ConvSPE when linear complexity is essential. Advanced methods demonstrate 5–6% gains.
6.5.2 Domain-Specific Guidelines
  • Biomedical signals: TUPE >> SPE >> T-PE for physiological complexity
  • Environmental sensors: SPE >> eRPE for regular sampling patterns
  • High-dimensional data ($d>5$): Advanced methods consistently outperform simple approaches
6.5.3 Computational Resource Framework
Limited resources:
Sinusoidal PE, tAPE ($\mathcal{O}(Ld)$ complexity)
Balanced scenarios:
SPE, TUPE (optimal accuracy–efficiency trade-off)
Performance-critical:
TUPE, SPE regardless of computational cost
6.5.4 Architecture-Specific Considerations
For Time Series Transformers, prioritize content-position separation methods (TUPE) and relative positioning approaches (eRPE, SPE). For Patch Embedding Transformers, multi-scale approaches (T-PE, ConvSPE) handle hierarchical processing more effectively.
6.6 Limitations & Future Directions
While this survey provides a broad benchmark, certain limitations should be acknowledged, which in turn highlight avenues for future research. First, our study, though diverse with eleven datasets, cannot encompass the full spectrum of time series characteristics, such as extremely high-frequency data, highly non-stationary series with abrupt concept drifts, irregularly sampled data, or datasets with extensive missing values. The performance of PE methods on such specialized data types remains an area for further investigation. Second, our architectural analysis was focused on two representative transformer variants. The interplay and relative performance of PE strategies might differ with other emerging or highly specialized transformer backbones (e.g., those designed for ultra-long sequences or specific forecasting tasks). Third, while we benchmarked eight prominent PE methods, the field is dynamic, and newer or more niche techniques might offer different advantages. Finally, although we employed a consistent training protocol, an exhaustive hyperparameter search for every PE method across all datasets and architectures was beyond the scope of this survey; further fine-tuning could potentially lead to marginal shifts in relative performance for specific combinations.
These limitations inform future work: Developing adaptive PEs that adjust to data characteristics is a key direction. Further research into hybrid PE strategies and integrating domain-specific temporal priors could yield significant advances. Deeper interpretability analysis of PE effects on attention mechanisms is also needed. Finally, extending this benchmark to other tasks (e.g., forecasting, anomaly detection) and a more detailed computational overhead analysis would be valuable.
7 Conclusion
In this survey, we have presented a comprehensive analysis of positional encoding methods in transformer-based time series models, examining their effectiveness across various time series tasks and data types. Through extensive experimentation with ten distinct encoding methods across eleven diverse datasets, we demonstrated that advanced methods like TUPE and SPE consistently outperform traditional approaches while maintaining computational efficiency. Our findings establish clear relationships between sequence characteristics and encoding method effectiveness, providing valuable guidance for practitioners in selecting appropriate methods based on their specific application requirements. These insights lay the foundation for future research in developing more efficient and adaptive positional encoding strategies for time series analysis.
Declarations
Conflicts of interest There are no conflicts of interest or competing interests.
Authors’ contributions
Habib Irani and Vangelis Metsis jointly conceptualized the study. Habib Irani conducted the literature review, implemented the experimental framework, and performed the benchmarking and analysis. Vangelis Metsis provided guidance on the methodology and contributed to the interpretation of results. Habib Irani wrote the initial manuscript draft, and Vangelis Metsis revised and edited the manuscript. All authors reviewed and approved the final version of the manuscript.